# On minimizing cyclists’ ascent times

### v3

@@ 1 Introduction
We begin this article by invoking a phenomenological model to estimate the power required to maintain a given speed. This model is analogous to the one used and discussed by Bos et al. (2023, 2024). Subsequently, we seek the riding strategy that results in the least ascent time, which — under stated constraints and assumptions — we show to be constant speed. We proceed to exemplify this result for several distinct ascents. We conclude by discussing the obtained result, comparing it to another plausible strategy: constant power, and commenting on possible future developments. In the appendices, we provide a method for imposing an optimization constraint, namely, the value of the available average power, justify an aspect of our model simplification, namely, neglecting the power used to increase kinetic energy, examine the effect of imposing the second constraint, namely, the value of the available maximum power, comment on numerical optimization, and discuss an ad hoc constant-ascent-speed strategy.
@@ 4.3 Time
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ Acknowledgements
We wish to acknowledge David Dalton, for his proofreading, Elena Patarini, for her graphic support, and Roberto Lauciello, for his artistic contribution.
@@ Appendix C Average-power constraint subject to maximum power
To gain insight to the effect of $P_{\rm max}$, let us restrict the power to $P_{\rm max}=400\,\rm W$. In accordance with expression (C.4), there are $3932$ extremal segments, along which $T_{P_{\rm max}}=116.3710\,\rm s$ and $W_{P_{\rm max}}=46548.4088\,\rm J$. For the submaximal segments, the constant speed is $V=4.3758\,\rm m/s$, which results in the ascent time of $T=254.9954\,\rm s$.
Refer to captionRefer to caption

### v5

@@ 1 Introduction
We begin this article by invoking a phenomenological model to estimate the power required to maintain a given speed. This model is analogous to the one used and discussed by Bos et al. (2023, 2024). Subsequently, we seek the riding strategy that results in the least ascent time, which — under stated constraints and assumptions — we show to be constant speed. We proceed to exemplify this result for several distinct ascents. We conclude by discussing the obtained result, comparing it to another plausible strategy: constant power, and commenting on possible future developments. In the appendices, we provide a method for imposing an optimization constraint, namely, the value of the available average power, justify an aspect of our model simplification, namely, neglecting the power used to increase kinetic energy, examine the effect of imposing the second constraint, namely, the value of the available maximum power, comment on numerical optimization, discuss an ad hoc constant-ascent-speed strategy, and conclude with an insight into empirical adequacy of the model.
@@ 5 Discussion and conclusions
Following recent measurements, in Appendix F, we examine briefly the empirical adequacy of our model. We do so for the constant-power strategy, which is easier to keep for a rider since it corresponds to steady effort.
@@ Acknowledgements
We wish to acknowledge Matteo Bertrand, for providing the measurements of his ascents, David Dalton, for his proofreading, Elena Patarini, for her graphic support, and Roberto Lauciello, for his artistic contribution.
@@ Appendix C Average-power constraint subject to maximum power
To gain insight to the effect of $P_{\rm max}$, let us restrict the power to $P_{\rm max}=400\,\rm W$. In accordance with expression (C.4), there are $3932$ extremal segments, along which $T_{P_{\rm max}}=116.3710\,\rm s$ and $W_{P_{\rm max}}=46548.4088\,\rm J$. For the submaximal segments, the constant speed is $V=4.3758\,\rm m/s$, which results in the ascent time of $T=255.3814\,\rm s$.
@@ Appendix E Constant ascent speed
Appendix F On empirical adequacy
The empirical adequacy of a phenomenological model needs to be examined by a comparison with measurements. For this process, we use the ascent — referred to as the San Bernardo segment — whose elevation gain is $162\,\rm m$. As shown in Figure F.1, the slope is variable; the average and maximum grades are $4.8\,\%$ and $8.4\,\%$, respectively. A recent ascent — performed specifically for our study — resulted in $T=543\,\rm s$, which is $9$ minutes and $3$ seconds.
Figure F.1: San Bernardo segment, which is a $3$rd category climb in the Italian foothills near Ovada; since the segment has three-dimensional coordinates, the horizontal axis is with respect to arclength, $s$.
The ascent was attempted with constant power, which is easier to maintain than constant speed, since it corresponds to a steady effort. As shown in Figure F.2, the measured power oscillates about its average of $\overline{P}=322.0699$, in such a manner that its standard deviation is $33.9902\,\rm W$. These oscillations are — at least in part — a consequence of cyclocomputer measurements which are sampled only once a second and, hence, occur necessarily at random instants of pedal revolutions.
Figure F.2: Comparison between average (black) and instantaneous (grey) power
To examine the model, we consider the following values in expression (1) : $m=78.6\,\rm kg$, as reported by the rider, ${\rm C_{d}A}=0.35\,{\rm m}^{2}$, ${\rm C_{rr}}=0.005$, $\lambda=0.02$, as estimated from standard measurements, $\rho=1.1464\,\rm kg/m^{3}$, as calculated based on altitude and temperature, and $g=9.81\,\rm m/s^{2}$. According to the constant-power strategy — with $P_{0}=322\,\rm W$ — $T=541.7489\,\rm s$, which is $9$ minutes and $2$ seconds; the ascent times agree to within one second.
Another agreement between the model and measurements is the instantaneous speed, shown in Figure F.3. This agreement is symptomatic of the model’s pertinence in examining certain subtleties, not only its capacity of predicting average or global quantities.
Figure F.3: Comparison between modelled (black) and measured (grey) instantaneous ground speeds
This initial examination supports the empirical adequacy of the model. More rides are planned in the foreseeable future to be followed by a statistical analysis. Notably, in accordance with our formulation, a constant-speed ascent — with $\overline{P}=322\,\rm W$ — should result in $T=527.4674\,\rm s$, which is $8$ minutes and $47$ seconds, a quarter of a minute less than a constant-power ascent.